# ATI-CTLO: Adaptive Temporal Interval-based Continuous-Time LiDAR-Only Odometry

### v2

@@ IV-A2 NTU VIRAL Dataset
The NTU VIRAL dataset records multiple sequences in several challenging indoor and outdoor conditions by an UAV, which equipped with two Ouster LiDAR: one horizontal (L1) and one vertical (L2). All algorithms are tested using the L1 LiDAR, and error evaluations are based on scripts provided by the dataset authors [^1] https://ntu-aris.github.io/ntu_viral_dataset/evaluation_tutorial.html. Since LeGO-LOAM is designed for ground robots, we employ another feature-based method F-LOAM instead. As shown in Table II, our algorithm achieve the highest accuracy in sequences eee, nya, sbs, rtp.
@@
The tnp sequences are recorded in a courtyard surrounded by three buildings, where the features in the z-axis direction are similar, and the horizontally placed L1 LiDAR provided fewer constraints on the z-axis compared to the x and y axes, exacerbating errors in the z-axis. In CT-ICP, each trajectory segment $[\mathbf{T}_{i},\mathbf{T}_{i+1}]$ is constrained by the corresponding complete LiDAR scan, offering more comprehensive constraints. Conversely, in our approach and Traj-LO, the segmentation of the point cloud results in significant vertical errors due to the similarity of local information in the horizontal direction. However, our approach to handling degeneracy effectively mitigates this negative impact, resulting in higher accuracy compared to Traj-LO. Fig. 6(a) shows the error components for each axis and environment of tnp_01.
In the spms sequences, the UAV flies away from the buildings, resulting in a lack of LiDAR points, causing the failure of most algorithms. Traj-LO exhibits robustness by ensuring odometry convergence through kinematic constraints. However, in the spms_02 sequence, it ultimately fails when the LiDAR points between control nodes become excessively sparse. In contrast, our method can detect degeneracy due to sparse LiDAR points and merges subsequent point cloud segments, ensuring a sufficient number of LiDAR points between control nodes, results are shown in Fig. 6(b).
@@ IV-A3 Our Own Jiulonghu Campus Dataset
In addition to public datasets, we validate the effectiveness of our method under challenge conditions. As shown in the Fig. 8, using the quadruped robot Jueying Lite2, equipped with a Robosense RS-Helios 32-line LiDAR, we collect challenging sequences on the lawn of Southeast University’s Jiulonghu Campus. The sequences labeled easy, medium, and hard increase in difficulty and distance from the lawn edge. In the easy sequence, the robot moves along the lawn’s edge. In the medium sequence, the robot moves towards the center of the lawn but still has some trees and other obstacles providing LiDAR points. In the hard sequence, the robot is in the middle of the lawn, surrounded only by the ground and sparse edge points, leading to significant degeneracy.
@@
easy
medium
hard
As shown in Table III, among the tested LiDAR-only odometry algorithms, ours is the only one that successfully completes all sequences. LeGO-LOAM and CT-ICP are unable to complete any test, while KISS-ICP manages to pass the easy sequences. Traj-LO manages well in easy and medium conditions but fails in hard sequence due to sparse feature points. Fig. 7 shows the result on hard sequence, CT-ICP experiences significant drift during aggressive turns, causing the program to terminate automatically. LeGO-LOAM drifts when moving towards the center of the lawn as the edge features gradually diminish. KISS-ICP and Traj-LO fail to handle changes in motion state when features are sparse. Our approach not only incorporates kinematic constraints but also dynamically increases temporal intervals during degeneracy, addressing the reduction of LiDAR information caused by cloud segmentation in continuous-time methods.
IV-B Real-Time Analysis
TABLE IV: The Pose Output Temporal Interval
@@ IV-B Real-Time Analysis
Our method dynamically adjusts the temporal interval between control nodes, resulting in a non-fixed pose output frequency. We evaluated the pose output frequency across sequences from three platforms. As depicted in the Table IV, the street_07 sequence is a standard road scene with continuous aggressive turns, yielding an average pose output rate of 64.16ms. In the spms_02 environment, degeneracy requires continuous cloud segment fusion, resulting in a lower output rate (82.89ms). In contrast, the hard sequence from the Jiulonghu Campus Dataset, despite continuous cloud segment fusion, has sparser environmental features leading to a higher output rate (48.78ms). The overall pose output rate of our method falls between 10Hz and 20Hz.
V CONCLUSIONS

### v3

@@ IV-A2 NTU VIRAL Dataset
The NTU VIRAL dataset, recorded by a UAV equipped with two Ouster LiDAR sensors (one horizontal, L1, and one vertical, L2), includes multiple sequences captured in various challenging indoor and outdoor conditions. All algorithms are tested using the L1 LiDAR, and error evaluations are based on script provided by the dataset authors [^1] https://ntu-aris.github.io/ntu_viral_dataset/evaluation_tutorial.html, which computes the RMSE of ATE. Since LeGO-LOAM is designed for ground robots, we employ another feature-based method F-LOAM [2] instead. As shown in Table II, our method achieves the highest accuracy in sequences eee, nya, sbs, rtp, spms. Notably, in the spms_02 sequence, the sparsity of environmental features (Fig. 6(b)) caused most algorithms to diverge. In contrast, our method, thanks to its degeneracy management, merges subsequent cloud segments and ensures stable performance. More details can be seen in Section. IV-B.
@@
In the tnp sequence, the surrounding buildings result in insufficient constraints along the z-axis (Fig. 6(a)). CT-ICP maintains lower ATE due to its broader neighborhood search range (27 closest voxels)[14]. Despite using a smaller search range of 7 closest voxels (similar to Traj-LO), our method achieves high accuracy, ranking second only to CT-ICP. The degeneracy management module helps by decreasing the downsampling rate ($D_{s}$) and merging point clouds when z-axis constraints are insufficient, minimizing the effect.
@@ IV-A3 Our Own Jiulonghu Campus Dataset
In addition to public datasets, we validate the effectiveness of our method under challenging conditions. As shown in Fig. 8, we collected challenging sequences using the quadruped robot Jueying Lite2, which is equipped with a Robosense RS-Helios 32-line LiDAR, on the lawn of Southeast University’s Jiulonghu Campus. The sequences labeled easy, medium, and hard increase in difficulty and distance from the lawn edge. In the easy sequence, the robot moves along the lawn’s edge. In the medium sequence, the robot moves towards the center of the lawn but still has some trees and other obstacles providing LiDAR points. In the hard sequence, the robot is in the middle of the lawn, surrounded only by the ground and sparse edge points, leading to significant degeneracy. As shown in Table III, among the tested LiDAR-only odometry algorithms, ours is the only one that successfully completes all sequences.
@@
easy(127.86m)
medium(92.44m)
hard(522.05m)
  • 1
    ’✗’ denotes significant error, as defined in the same way as in Table I
Notably, in the hard sequence, similar to the spms_02 sequence from the NTU dataset, our method shows remarkable performance in environments with sparse features. Results are shown in Fig. 7.
IV-B Ablation Studies
TABLE IV: RMSE of ATE (m) on Ablation Studies
Approach Newer College Dataset NTU VIRAL Dataset
05_Quad(479.0m) 06_Dynamic(97.2m) 07_Parkland(695.7m) spms_02(697.9m)
KISS-ICP 0.298 0.260
CT-ICP 0.098 0.143 0.180
Traj-LO 0.334 0.146
Ours(None) 0.186
Ours(PCA) 0.124 0.131
Ours(DM) 0.175 0.184 0.126
Ours(PCA+DM) 0.119 0.124 0.130 0.077
  • 1
    ’✗’ denotes significant error, as defined in the same way as in Table I.  ’(*m)’ denotes the length of the sequence.
  • 2
    The results of CT-ICP on Newer College Dataset are obtained from[13].
To further validate the effectiveness of our method, ablation studies are conducted on the Newer College Dataset, with the results presented in Table IV. Ours(None) represents our proposed method with both the PCA pre-assessment module and degeneracy management disabled. Ours(PCA) and Ours(DM) refer to Ours(None) with the addition of either PCA pre-assessment or the Degeneracy Management, respectively. Ours(PCA+DM) represents the full implementation of our proposed method.
The importance of the PCA pre-assessment in dynamic motion is demonstrated in the 06_Dynamic and 07_Parkland sequences. In the 06_Dynamic sequence, characterized by aggressive motion with rotational speeds up to 3.5 rad/s, the addition of PCA results in a 13% accuracy improvement compared to the SOTA methods. In 07_ParkLand, sudden motion changes cause failures in Ours(None) and Ours(DM), but the PCA module reduces the temporal interval, improving the algorithm’s effectiveness.
Ablation experiments on the spms_02 and 05_Quad sequences highlight the robustness enhancement brought by the degeneracy management module. In the 05_Quad sequence, Ours(None) and Ours(PCA) diverge due to the narrow starting space, while Ours(DM) adapts to the indoor environment with degeneracy management. Ours(PCA+DM) further improves outdoor accuracy with PCA pre-assessment. In the spms_02 sequence, Ours(DM) integrates subsequent point cloud data to acquire sufficient observational constraints, ensuring the stability of the algorithm.
IV-C Real-Time Analysis
TABLE V: The Pose Output Temporal Interval
@@ IV-C Real-Time Analysis
Our method dynamically adjusts the temporal interval between control nodes, resulting in a non-fixed pose output frequency. We evaluated the pose output frequency across sequences from three platforms. As depicted in the Table V, the street_07 sequence is a standard road scene with continuous aggressive turns, yielding an average pose output rate of 64.16ms. In the spms_02 environment, degeneracy requires continuous cloud segment fusion, resulting in a lower output rate (82.89ms). In contrast, the hard sequence from the Jiulonghu Campus Dataset, despite continuous cloud segment fusion, has sparser environmental features leading to a higher output rate (48.78ms). The overall pose output rate of our method falls between 10Hz and 20Hz.
V Conclusion